# Slot-BERT: Self-supervised Object Discovery in Surgical Video

### v3

@@ Abstract
Object-centric slot attention is a powerful framework for unsupervised learning of structured and explainable representations that can support reasoning about objects and actions, including in surgical video. While conventional object-centric methods for videos leverage recurrent processing to achieve efficiency, they often struggle with maintaining long-range temporal coherence required for long videos in surgical applications. On the other hand, fully parallel processing of entire videos enhances temporal consistency but introduces significant computational overhead, making it impractical for implementation on hardware in medical facilities. We present Slot-BERT, a bidirectional long-range model that learns object-centric representations in a latent space while ensuring robust temporal coherence. Slot-BERT scales object discovery seamlessly to long videos of unconstrained lengths. A novel slot contrastive loss further reduces redundancy and improves the representation disentanglement by enhancing slot orthogonality. We evaluate Slot-BERT on real-world surgical video datasets from abdominal, cholecystectomy, and thoracic procedures. Our method surpasses state-of-the-art object-centric approaches under unsupervised training achieving superior performance across diverse domains. We also demonstrate efficient zero-shot domain adaptation to data from diverse surgical specialties and databases.
@@ 1 Introduction
These issues become especially pronounced in learning disentangled representation for scene decomposition in surgical videos, where temporal dynamics can be highly complex. Instruments and tissues in surgical videos often move at different speeds, and the visibility of specific objects can fluctuate significantly over the course of a procedure. Moreover, many current object-centric learning approaches for videos fail to fully exploit the sequential structure of longer episodes. In surgical videos, a typical episode depicting a surgical action or a surgical task will last from a few seconds to a minute or longer. RNN-based models (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022) struggle with learning from sequences of this length. Employing slot-attention mechanisms to process entire video sequences in parallel, as proposed in (Singh et al., 2024), could help overcome some challenges but may struggle with scalability for longer videos in real-world datasets due to computational limitations.
@@ 2.1 Self-supervised object-centric learning
Learning of object-centric representations can demonstrably improve the sample efficiency and generalization of vision and dynamics models in compositional domains. In order to discover objects from images or videos without supervision, the learning architecture must bind the distinct features that belong to particular objects and create representations with activation patterns that are distinct for object instances. Various approaches have been proposed to achieve this goal. Early works focused on variational auto-encoders (VAEs) (Kingma & Welling, 2014) and disentangled representation learning (Mathieu et al., 2019; Eastwood & Williams, 2018; Kim & Mnih, 2018; Higgins et al., 2017). Other efforts employed iterative attention mechanisms, such as Capsule Networks (Sabour et al., 2017; Hinton et al., 2018; Tsai et al., 2020), contrastive learning approaches (Kipf et al., 2020; Hénaff et al., 2022; Xu et al., 2022), or complex-valued auto-encoders that implicitly encode temporal correlation (Löwe et al., 2022). Beyond learning generalized representations, reconstructive and generative object-centric models also show the ability to segment objects in scenes (Lin et al., 2020; Van Steenkiste et al., 2020; Greff et al., 2019; Engelcke et al., 2020; Burgess et al., 2019).
@@ 4.1 Dataset
Cholec80: The Cholec80 dataset (Twinanda et al., 2016) contains 80 cholecystectomy surgery videos downsampled to 1 FPS with presence labels for seven surgical tools in each frame. Following the preprocessing steps of a weakly supervised learning study (Liao et al., 2024), we derived a Transient Object Presence (TOP) video subset from Cholec80 resulting in 5,296 clips with 30 frames each, in which objects are not necessarily present in every frame, but can rather transition in or out of the field of view. The test set is based on 100 clips from the CholecSeg8K dataset (Hong et al., 2020), which contains 8,000 frames with instrument and anatomy segmentation from 17 Cholec80 videos.
@@
To demonstrate the benefit of applying our method to broader domains, We also evaluate Slot-BERT on non-surgical datasets, including a natural real-world video dataset YT-VIS (Xu et al., 2019) and a synthetic multi-object tracking data MOVi-E (Greff et al., 2022). Results on these two datasets are concluded in Supplementary material F.
@@ 4.2 Metrics
We evaluate our approach based on the quality of the slot masks produced by the decoder using four primary metrics: video foreground ARI (FG-ARI) (Greff et al., 2019), video mean best overlap (mBO) (Pont-Tuset et al., 2016), mean best Hausdorff Distance (mBHD) and CorLoc(Bilen & Vedaldi, 2016) for evaluating localization accuracy. FG-ARI, adapted from a widely-used metric in object-centric research, measures the similarity between predicted object masks and ground truth masks, focusing on how effectively objects are segmented. mBO, on the other hand, evaluates the alignment between predicted and ground truth masks using the intersection-over-union (IoU) metric. For mBO, each ground truth mask is matched to the predicted mask with the highest IoU via Hungarian matching, and the average IoU is computed across all matched pairs. While FG-ARI primarily measures segmentation quality, mBO provides a more comprehensive assessment by including background pixels. Furthermore, the video version of mBO (mBO-V) also accounts for the temporal consistency of masks throughout the video. To facilitate comparisons at the frame level, we include the image-based version of mBO (mBO-F), which is computed on individual frames. This metric evaluates segmentation quality at the image level without considering temporal consistency. Additionally, to assess the boundary accuracy of the predicted masks, we calculate the mean best Hausdorff Distance (mBHD) for the best-overlapping masks produced by the object-centric learning model. CorLoc is also calculated based on overlapped object instance bounding box, given a threshould IoU>>0.5.
@@ 4.4 Experiment setup
All models were trained with the first 5 frames of each clip, even when the clips had longer sequences. For testing, each video was repeatedly processed by the trained model three times to calculate an average score and standard deviation, ensuring that any potential instabilities in the model’s outputs were accounted for (Wu et al., 2023b).
Finally, we tested performance on longer frame sequences of 7, 11, and 30 frames and evaluated a variation of our model on longer sequences by duplicating the TST module as a next-slot initializer.
@@ 4.7 Zero-shot performance
4.8 Adaptation to longer sequences
@@ 7 Frames 11 Frames
To evaluate the robustness and generalization of our method to longer video sequences, we tested our model on sequences of 7 and 11 frames after training on 5 frames. A sliding window approach was employed to handle longer inputs, ensuring temporal alignment during prediction. Additionally, we introduced a future slot prediction mechanism that replaced the simple RNN-based slot initialization which is only possible when testing on video length that is equal to length in training. In this new design, the temporal slot transformer is fed all previous T1T-1 slots, leaving the slot at the last position empty. The temporal slot transformer then predicts this missing slot, which serves as the initialization for subsequent frames. An illustration of this next slot initialization can be found in supplementary material section A.
@@
Table 7: Performance comparison of different model configurations in the ablation study. The results highlight the contributions of individual components, including the contrastive loss, slot-BERT module, and slot-specific masking, under both same-domain (MICCAI) and zero-shot transfer scenarios (EndoVis and Thoracic datasets).
Same domain (MICCAI) Zero shot transfer
5 frames 11 frames Endovis Thoracic
w/o contrast w/o TST 45.0 ± 0.5 53.5 ± 0.5 42.0 ± 0.5 53.2 ± 0.6 39.5 ± 0.2 50.1 ± 0.2 34.5 ± 0.1 36.6 ± 0.1
contrast only 46.4 ± 0.6 55.3 ± 0.6 44.1 ± 0.1 55.0 ± 0.2 42.1 ± 0.4 52.6 ± 0.3 34.2 ± 0.1 37.5 ± 0.2
w/o contrast 46.3 ± 0.4 55.2 ± 0.4 43.7 ± 0.3 54.7 ± 0.1 40.8 ± 0.3 51.5 ± 0.2 34.1 ± 0.2 36.7 ± 0.2
@@ 4.9 Ablation study
To investigate the contributions of individual components in our model to its final performance, we systematically removed or modified different components and observed their impact on segmentation accuracy, as summarized in Table 7. These experiments were conducted in both same-domain and zero-shot transfer scenarios, with evaluations on the MICCAI dataset (short and long sequences), EndoVis dataset, and Thoracic dataset. Below, we discuss the implications of the results for each configuration.
Baseline: In this configuration, both the contrastive loss and TST module were removed, and a simple recurrent refinement network (RRN) was employed along with a standard four-layer MLP decoder (denoted as w/o contrast w/o TST). This baseline yielded the lowest performance across all datasets and settings. For instance, on the MICCAI test set (long sequences), the mBO-V score dropped to $42.0\%$, while on EndoVis, it fell to $39.5\%$. These results highlight the critical contributions of both the contrastive loss and the TST module in enhancing segmentation quality.
Slot Contrast only: Removing the slot-BERT module while retaining the contrastive loss (denoted as contrast only) resulted in improvements over the vanilla baseline. For instance, on the MICCAI dataset (short sequences), the mBO-V score increased to $46.4\%$, and the FG-ARI improved to $55.3\%$. However, the performance remained below that of the full model, underscoring the role of slot-BERT in refining slot representations and contextualizing features effectively.
Effect of slot contrastive loss: In this configuration, the contrastive loss was omitted while retaining the slot-BERT module (denoted as w/o contrast). The results were similar to those of the previous configuration, with slight decreases in performance on most datasets. For example, on the EndoVis dataset, the mBO-V score was $40.8\%$, which is $1.3\%$ lower than the configuration without TST but with contrastive loss. These findings indicate that the contrastive loss is pivotal for improving slot distinctiveness, particularly in datasets with diverse instrument appearances.
Effect of slot masks: We also examined the impact of slot-specific masking by directly inputting video slots into the TST module without applying slot masks and replacing the masked transformer with a standard transformer (referred to as w/o slot masks). This led to a slight performance degradation. For instance, on the MICCAI dataset (long sequences), the mBO-V score was $45.9\%$ compared to $46.9\%$ for the full model. This suggests that slot masks do contribute to robust representation learning by focusing on relevant feature regions.
The results in Table 7 demonstrate the critical role of each component in our model. The full model consistently outperformed all other configurations across datasets and metrics. Notably, it achieved the highest mBO-V scores in all settings, including $48.9\%$ on MICCAI (short sequences), $44.0\%$ on EndoVis, and $37.7\%$ on Thoracic. These results validate the synergistic contributions of the TST module, contrastive loss, and slot-specific masking in achieving superior segmentation performance. Specifically, the contrastive loss enhances inter-slot separation, the TST module improves feature contextualization and representation, and slot-specific masking ensures better focus on relevant regions. The full model’s superior performance across both same-domain and zero-shot scenarios underscores the importance of integrating these components for robust segmentation in diverse domains.
Figure 5: Slot masks and the corresponding cosine similarity matrices for Slot-BERT and ablated Slot-BERT variants without the TST module (left) or contrastive loss (right). All models are trained with 7 slots, and each segmentation mask is annotated with its corresponding slot index (0–6). See text for details.
@@ 4.10 Contrastive hyperparameter tuning analysis
Figure 6: Parameter tuning of Slot-BERT with respect to contrastive loss weight ($\alpha$) and temperature ($\tau$). Performance is reported as mBO-V. Across a range of $\alpha$ values (0.0001–0.1) and $\tau$ values (0.05–100), we observe consistent improvement compared to the baseline without contrastive learning. The best performance is achieved with $\alpha$ = 0.01 and $\tau$ = 0.5, reaching 50.1 mBO-V, highlighting the importance of jointly balancing temperature scaling and loss weighting.
@@ 4.11 Experiment with slot-Mixer Decoder
These findings confirm that our method is able to surpass the vanilla RNN-based baseline using either MLP broadcast, or Slot-Mixer decoder which improves decoding quality by introducing interactions between slots, achieving high localization and segmentation performance. For more qualitative results refer to the section G.1 of supplementary materials.
@@ 4.13 Experiment with different slot numbers
In order to investigate the effect of changing slot number $K$ on the performance of Slot-BERT, experiments were conducted using MLP and Mixer decoders. Figure 8 presents the CorLoc score obtained on training domain and zeroshot data, to evaluate impact of slot allocation on the localization accuracy. The results show that $K=7$ generally provides the best performance for both methods, with MLP achieving a peak score of 70.7 and Mixer reaching 67.4 on training domain data. As $K$ increases beyond 7, performance begins to decline for both decoders, suggesting that excessively large $K$ values might over-delineate the image as the number of objects in a video is limited. MLP consistently outperforms Mixer on training domain data, indicating its superior ability to leverage increased $K$ for improved predictions.
On zero-shot data, the trend remains consistent, with both MLP and Mixer achieving their best scores at $K=7$, though their overall performance is lower compared to the training domain. MLP slightly outperforms Mixer across most KKvalues, reaching a maximum zero-shot score of 62.8, compared to Mixer’s 61.9 at $K=7$. These results highlight that while $K=7$ is optimal across both datasets, there is a clear performance gap between training domain and zeroshot data, reflecting challenges in generalization. The analysis demonstrates that adjusting $K$ significantly impacts the effectiveness of both decoding methods, emphasizing the importance of finding an optimal KKvalue to balance performance and generalization. Evaluation of different slot number under additional metrics are concluded in Supplementary Material section D.

### v4

# arXiv:2501.12477v4[eess.IV] 02 Mar 2026

@@ Abstract
Object-centric slot attention is a powerful framework for unsupervised learning of structured and explainable representations that can support reasoning about objects and actions, including in surgical video. While conventional object-centric methods for videos leverage recurrent processing to achieve efficiency, they often struggle with maintaining long-range temporal coherence required for long videos in surgical applications. On the other hand, fully parallel processing of entire videos enhances temporal consistency but introduces significant computational overhead, making it impractical for implementation on hardware in medical facilities. We present Slot-BERT, a bidirectional long-range model that learns object-centric representations in a latent space while ensuring robust temporal coherence. Slot-BERT scales object discovery seamlessly to long videos of unconstrained lengths. A novel slot contrastive loss further reduces redundancy and improves the representation disentanglement by enhancing slot orthogonality. We evaluate Slot-BERT on real-world surgical video datasets from abdominal, cholecystectomy, and thoracic procedures. Our method surpasses state-of-the-art object-centric approaches under unsupervised training achieving superior performance across diverse domains. We also demonstrate efficient zero-shot domain adaptation to data from diverse surgical specialties and databases. Code and data shared at: https://github.com/PCASOlab/slot-BERT.
@@ 1 Introduction
These issues become especially pronounced in learning disentangled representation for scene decomposition in surgical videos, where temporal dynamics can be highly complex. Instruments and tissues in surgical videos often move at different speeds, and the visibility of specific objects can fluctuate significantly over the course of a procedure. Moreover, many current object-centric learning approaches for videos fail to fully exploit the sequential structure of longer episodes. In surgical videos, a typical episode depicting a surgical action or a surgical task will last from a few seconds to a minute or longer. (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022) struggle with learning from sequences of this length. Employing slot-attention mechanisms to process entire video sequences in parallel, as proposed in (Singh et al., 2024), could help overcome some challenges but may struggle with scalability for longer videos in real-world datasets due to computational limitations.
@@ 2.1 Self-supervised object-centric learning
Learning of object-centric representations can demonstrably improve the sample efficiency and generalization of vision and dynamics models in compositional domains. In order to discover objects from images or videos without supervision, the learning architecture must bind the distinct features that belong to particular objects and create representations with activation patterns that are distinct for object instances. Various approaches have been proposed to achieve this goal. Early works focused on variational auto-encoders (VAEs) (Kingma and Welling, 2014) and disentangled representation learning (Mathieu et al., 2019; Eastwood and Williams, 2018; Kim and Mnih, 2018; Higgins et al., 2017). Other efforts employed iterative attention mechanisms, such as Capsule Networks (Sabour et al., 2017; Hinton et al., 2018; Tsai et al., 2020), contrastive learning approaches (Kipf et al., 2020; Hénaff et al., 2022; Xu et al., 2022), or complex-valued auto-encoders that implicitly encode temporal correlation (Löwe et al., 2022). Beyond learning generalized representations, reconstructive and generative object-centric models also show the ability to segment objects in scenes (Lin et al., 2020; Van Steenkiste et al., 2020; Greff et al., 2019; Engelcke et al., 2020; Burgess et al., 2019).
@@ 4.1 Dataset
Cholec80: The Cholec80 dataset (Twinanda et al., 2016) contains 80 cholecystectomy surgery videos downsampled to 1 FPS with presence labels for seven surgical tools in each frame. Following the preprocessing steps of a weakly supervised learning study (Liao et al., 2024), we derived a Transient Object Presence (TOP) video subset from Cholec80 resulting in 5,296 clips with 30 frames each, in which objects are not necessarily present in every frame, but can rather transition in or out of the field of view. The test set is based on 100 clips from the CholecSeg8K dataset (Hong et al., 2020), which contains 8,000 frames with instrument and anatomy segmentation from 17 Cholec80 videos. To evaluate long-range object-centric tracking under challenging multi-instrument interaction conditions, we curate a separate test set of 20 unseen cholecystectomy videos, each comprising 30-frame (30 s) sequences. These videos feature frequent instrument interactions, occlusions, and repeated instrument entry and exit events, posing a substantial challenge for maintaining temporal identity consistency.
@@
To demonstrate the benefit of applying our method to broader domains, We also evaluate Slot-BERT on non-surgical datasets, including a natural real-world video dataset YT-VIS (Xu et al., 2019) and a synthetic multi-object tracking data MOVi-E (Greff et al., 2022). Results on these two datasets are concluded in Supplementary material H.
@@ 4.2 Metrics
We evaluate our approach based on the quality of the slot masks produced by the decoder using four primary metrics: video foreground ARI (FG-ARI) (Greff et al., 2019), video mean best overlap (mBO) (Pont-Tuset et al., 2016), mean best Hausdorff Distance (mBHD) and CorLoc(Bilen and Vedaldi, 2016) for evaluating localization accuracy. FG-ARI, adapted from a widely-used metric in object-centric research, measures the similarity between predicted object masks and ground truth masks, focusing on how effectively objects are segmented. mBO, on the other hand, evaluates the alignment between predicted and ground truth masks using the intersection-over-union (IoU) metric. For mBO, each ground truth mask is matched to the predicted mask with the highest IoU via Hungarian matching, and the average IoU is computed across all matched pairs. While FG-ARI primarily measures segmentation quality, mBO provides a more comprehensive assessment by including background pixels. Furthermore, the video version of mBO (mBO-V) also accounts for the temporal consistency of masks throughout the video. To explicitly evaluate long-range temporal consistency, we additionally report identity-aware tracking metrics, including the slot-level IDF1 score (Ristani et al., 2016) and Temporal Identity Persistence (T-IDP) (Milan et al., 2016). IDF1 measures identity precision and recall across frames, while T-IDP quantifies the fraction of frames in which each ground-truth instance is consistently associated with the same predicted slot. To facilitate comparisons at the frame level, we include the image-based version of mBO (mBO-F), which is computed on individual frames. This metric evaluates segmentation quality at the image level without considering temporal consistency. Additionally, to assess the boundary accuracy of the predicted masks, we calculate the mean best Hausdorff Distance (mBHD) for the best-overlapping masks produced by the object-centric learning model. CorLoc is also calculated based on overlapped object instance bounding box, given a threshould IoU>>0.5.
@@ 4.4 Experiment setup
For testing, videos were processed by the trained model three times to calculate an average score and standard deviation, ensuring that any potential instabilities in the model’s outputs were accounted for (Wu et al., 2023b). Five processing repetitions were used in ablation studies.
For testing sequences longer than the training context window, we adopt an online recurrent inference strategy with a sliding window. In addition, we evaluate a variant of our model that extends temporal reasoning to longer sequences by duplicating the TST module and using it as a next-slot initializer.
@@ 4.7 Zero-shot performance
4.8 Longer and more challenging sequences
@@
Figure 5: Box plot comparison of Slot-BERT and the baseline in a challenging scenario characterized by frequent object entry and exit. Long-range temporal tracking performance is assessed using IDF1 and Temporal Identity Persistence (T-IDP), together with object localization accuracy (CorLoc).
@@ 7 Frames 11 Frames
To evaluate the robustness and generalization of our method to longer video sequences, we tested our model on sequences of 7 and 11 frames after training on 5 frames. A sliding window approach was employed to handle longer inputs, ensuring temporal alignment during prediction. Additionally, we introduced a future slot prediction mechanism that replaced the simple RNN-based slot initialization which is only possible when testing on video length that is equal to length in training. In this new design, the temporal slot transformer is fed all previous T1T-1 slots, leaving the slot at the last position empty. The temporal slot transformer then predicts this missing slot, which serves as the initialization for subsequent frames. An illustration of this next slot initialization can be found in supplementary material section B.
@@
We evaluate long-range temporal tracking under more challenging conditions, using the additional test set of unseen Cholecystectomy videos, where frequent object occlusions as well as instrument entry and exit occur. This setting reflects realistic surgical workflows and poses a significant challenge for online object-centric tracking. Long-range tracking performance is assessed using identity-aware metrics, including IDF1 and Temporal Identity Persistence (T-IDP), together with object localization accuracy (CorLoc).
As shown in Fig. 5, we evaluate tracking over sequences of 30 frames and compare our full model against a baseline using vanilla RNN-based inference without TST or contrastive learning. Our real-time method improves average slot-level IDF1 by 9.4%, T-IDP by 8.9%, and CorLoc by 8.8%, demonstrating substantially stronger identity consistency and localization accuracy under frequent object entry and exit.
@@ 4.9 Ablation study
We conduct ablation studying by isolating the effects of our two main contributions, namely the slot contrastive objective and the TST module, on video segmentation accuracy and temporal tracking consistency across different datasets and video lengths.
All ablation experiments are repeated across five independent runs, resulting in 500-1000 total video-level evaluations per experiment. Statistical significance is assessed using one-tailed Welch’s $t$-tests and Bonferroni correction for multiple comparisons. The observed improvements between the full model and single-component ablations correspond to large effect sizes (Cohen’s $d>0.8$), indicating that the gains are practically meaningful.
@@
Figure 7 presents ablation results on three datasets with varying test-video lengths. We evaluate video-level accuracy (mBO-V) and temporal-slot IDF1, comparing the full method against variants with (w) or without (w/o) the contrastive objective and/or the TST module. Note that the Thoracic dataset is only partially and sparsely annotated; therefore, comparisons on this dataset are conducted only at full video length (30 frames).
Across all datasets, both contrastive learning and temporal self-training consistently improve surgical instrument segmentation and tracking. On the MICCAI dataset (11-frame sequences), adding contrastive learning alone improves mBO-V by approximately 3.6 percentage points over the baseline, while adding TST alone yields an improvement of approximately 5.1 points; combining both components results in a larger gain of approximately 6.7 points. Similar trends are observed across datasets and metrics, with combined gains of 3–7 points in mBO-V and 4–11 points in temporal slot IDF1.
Most of these improvements are statistically significant ($p<0.0167$, with Bonferroni correction $m=3$). An exception is observed for mBO-V on Cholec at 30 frames when comparing the full model to the TST-only variant, where the difference does not reach statistical significance ($p=0.46$). This behavior likely reflects the increased difficulty of segmenting tissue and small instruments in Cholec videos. Importantly, even in this setting, the full model achieves significantly higher temporal slot IDF1 than the TST-only variant ($p<0.01$), indicating stronger identity consistency. Taken together, these results confirm the complementary roles of contrastive learning and temporal self-training in robust surgical scene decomposition and long-range tracking.
Figure 6: Slot masks and the corresponding cosine similarity matrices for Slot-BERT and ablated Slot-BERT variants without the TST module (left) or contrastive loss (right). All models are trained with 7 slots, and each segmentation mask is annotated with its corresponding slot index (0-6). See text for details.
Figure 7: Ablation study on three datasets with varying test video lengths. We compare video-level segmentation accuracy (mBO-V) and the temporal slot IDF1 score. Our full method is evaluated against variants with (w) or without (w/o) the slot contrastive learning objective and the TST module. Note that the Thoracic dataset is only partially and sparsely annotated; therefore, comparisons on this dataset are conducted only using the full video length (30 frames).
@@ 4.10 Contrastive hyperparameter tuning analysis
Figure 8: Parameter tuning of Slot-BERT with respect to contrastive loss weight ($\alpha$) and temperature ($\tau$). Performance is reported as mBO-V. Across a range of $\alpha$ values (0.0001-0.1) and $\tau$ values (0.05-100), we observe consistent improvement compared to the baseline without contrastive learning. The best performance is achieved with $\alpha$ = 0.01 and $\tau$ = 0.5, reaching 50.1 mBO-V, highlighting the importance of jointly balancing temperature scaling and loss weighting.
@@ 4.11 Experiment with Slot-Mixer decoder
These findings confirm that our method is able to surpass the vanilla RNN-based baseline using either MLP broadcast, or Slot-Mixer decoder which improves decoding quality by introducing interactions between slots, achieving high localization and segmentation performance. For more qualitative results refer to the section I.1 of supplementary materials.
@@ 4.13 Experiment with different slot numbers
In order to investigate the effect of changing slot number $K$ on the performance of Slot-BERT, experiments were conducted using MLP and Mixer decoders. Figure 10 presents the CorLoc score obtained on training domain and zero-shot data, to evaluate impact of slot allocation on the localization accuracy. The results show that $K=7$ generally provides the best performance for both methods, with MLP achieving a peak score of 70.7 and Mixer reaching 67.4 on training domain data. As $K$ increases beyond 7, performance begins to decline for both decoders, suggesting that excessively large $K$ values might over-delineate the image as the number of objects in a video is limited. MLP consistently outperforms Mixer on training domain data, indicating its superior ability to leverage increased $K$ for improved predictions.
On zero-shot data, the trend remains consistent, with both MLP and Mixer achieving their best scores at $K=7$, though their overall performance is lower compared to the training domain. MLP slightly outperforms Mixer across most KKvalues, reaching a maximum zero-shot score of 62.8, compared to Mixer’s 61.9 at $K=7$. These results highlight that while $K=7$ is optimal across both datasets, there is a clear performance gap between training domain and zeroshot data, reflecting challenges in generalization. The analysis demonstrates that adjusting $K$ significantly impacts the effectiveness of both decoding methods, emphasizing the importance of finding an optimal KKvalue to balance performance and generalization. Evaluation of different slot number under additional metrics are concluded in Supplementary Material section E.